In a cars carrier, how many ways are there to put 8 cars labeled as {A, B, ..., H} in 4 containers labeled as { 1, 2, 3, 4 }? Each container must include at least one car.
Your answer is
40824

If the containe #1 must have the car A in it, find the new number of possible ways.
Your answer is
5040
are my answers correct?

Found 2 solutions by ikleyn, Edwin McCravy:
Answer by ikleyn(52946)

.

For the first question, it is not YOUR answer.

It is MY answer, which I provided to you in my post

https://www.algebra.com/algebra/homework/Probability-and-statistics/Probability-and-statistics.faq.question.1205233.html


Call the things by their proper names.



Answer by Edwin McCravy(20067)

I'll just do the first part. There are just these 5 ways to
break up 8 cars into a sum of 4 positive integers, without 
regard for order:
1. 1+1+1+5 = 8
2. 1+1+2+4 = 8
3. 1+1+3+3 = 8
4. 1+2+2+3 = 8
5. 2+2+2+2 = 8

Of course we must choose and order each of the 5 cases.

Case 1: 5 cars in 1 container, 1 car in each of the other 3. 
We can choose the container to hold 5 cars in 4 ways. 
We can choose the 5 cars for that container C(8,5)=56 ways. The other
3 cars can be distributed to the other 3 containers in 3!=6 ways.
That's (4)(56)(6) = 1344 ways  

Case 2: 4 cars in 1 container, 2 cars in another container, and 1 car in
each of the other 2 containers. 
We can choose the container to hold 4 cars in 4 ways.
We can choose the 4 cars for that container C(8,4)=70 ways.
We can choose the container to hold 2 cars in 3 ways.
We can choose the 2 cars for that container in C(4,2)=6 ways. The other
2 cars can be distributed to the other 2 containers in 2!=2 ways.
That's (4)(70)(6)(2) = 3360 ways

Case 3: 3 cars each in 2 containers, and 1 car in each of the other 2 containers. 
We can choose the 2 containers to hold 3 cars in C(4,2)=6 ways.
We can choose the 3 cars for the container with the smaller number C(8,3)=56 ways.
We can choose the cars for the container with the larger number C(5,3)=10 ways.
The remaining 2 cars can be distributed to the other 2 containers in 2!=2 ways.
That's (6)(56)(10)(2) = 6720 ways  

Case 4: 2 cars each in 2 containers, 3 cars in 1 container, and 1 car in 1 container  
We can choose the 2 containers to hold 2 cars in C(4,2)=6 ways.
We can choose the 2 cars for the container with the smaller number C(8,2)=28 ways.
We can choose the 2 cars for the container with the larger number C(6,2)=15 ways.
We can choose the container to hold 3 cars in C(2,1)=2 ways.
We can choose the 3 cars for that container in C(4,3)=4 ways.
The container to hold only 1 car gets the 1 car left in 1 way.
That's (6)(28)(15)(2)(4)(1)=20160 ways 

Case 5: Each container holds 2 cars each.
We can choose the 2 cars for the container with the number 1 in C(8,2)=28
ways.
We can choose the 2 cars for the container with the number 2 in C(6,2)=15 ways.
We can choose the 2 cars for the container with the number 3 in C(4,2)=6 ways.
We can choose the 2 cars for the container with the number 4 in C(2,2)=1 way.
That's (28)(15)(6)(1)=2520

The total = 1344+3360+6720+20160+2520 = 34104

I haven't gone over this with a fine-tooth comb to see if I've made an error.
I'll check it later, and correct it, or maybe another tutor will solve it. 

Edwin